The length of a rectangle is seven feet more than its width if the perimete is 18 feet what are its dimensions
Answer by josgarithmetic(39799)

L=w+7, 2w+2L=18.

L-w=7 and 2L+2w=18
OR
L-w=7 and L+w=9. Now this can be used right-away in the Elimination Method.